OpenStudy (anonymous):

y=x^2-6x+2 . Explain how to identidy the vertex and two other points on the parabola. I also need to sketch it but i have no clue how

OpenStudy (abb0t):

Start by completeing the square.

OpenStudy (anonymous):

the quadrent?

OpenStudy (abb0t):

Use: \[-\frac{ b }{ 2a }\]

OpenStudy (abb0t):

then, plug it in to the quadratic.

OpenStudy (anonymous):

i got -3

OpenStudy (abb0t):

Now, plug it back in to get your y value.

OpenStudy (abb0t):

f(-3)

OpenStudy (anonymous):

so y=3^2-6(3)+2 -7?

OpenStudy (abb0t):

Did you plug in 3 or (-3)?

OpenStudy (anonymous):

oops 29 i did 3

OpenStudy (raden):

it should be f(3), so recalculate the value of f(3) = ?

OpenStudy (anonymous):

I got 29

OpenStudy (raden):

-b/2a = -(-6)/2*1 = 6/2 = 3 right ?

OpenStudy (anonymous):

lol yeah, I missed yeah the equation above

OpenStudy (raden):

now, calculating again the value of f(3) = 3^2 - 6(3) + 2 = 9 - 18 + 2 = ?

OpenStudy (anonymous):

the answer is -7?

OpenStudy (raden):

correct

OpenStudy (anonymous):

so thats my vertex?

OpenStudy (raden):

so, the vertex is (3, -7)

OpenStudy (anonymous):

ohhhh YAE!

OpenStudy (anonymous):

How would you find 2 other pionts, just plug a # into the equation?

OpenStudy (raden):

what do u meant " two other points" ? is it the x-intercept and y-intercept ?

OpenStudy (anonymous):

I think it wants me to find one point and the mirror image if it.

OpenStudy (raden):

use wolfram :), ur sketch here : http://www.wolframalpha.com/input/?i=y%3Dx^2-6x%2B2

OpenStudy (raden):

but if u want do by manual, u have to determine the values of x-intercept and y-intercept also